OpenStudy (anonymous):

A flatbed truck can haul 60 crates in 5 loads. How many crates can it haul in 8 loads? A. 300 B. 480 C. 96 D. 68

OpenStudy (ayubie):

60/5 = x/8 8(60/5) = x x = 96 C.